OpenStudy (anonymous):

Identify the missing factors: 20 = (4) ( )

OpenStudy (anonymous):

5

OpenStudy (anonymous):

lets pretend the missing factor = x 20 = (4) (x) so, x = 20/4 x = 5 understand? :)